A company is projected to generate free cash flows of $47 million per year for the next two years, after which it is projected grow at a steady rate in perpetuity. The company's cost of capital is 11.2%. It has $24 million worth of debt and $6 million of cash. There are 14 million shares outstanding. If the appropriate terminal exit value for this company is 14, what's your estimate of the company's stock price? Round to one decimal place.
Answer #1

FCF in year 1 = $ 47 M

FCF in year 2 = 47 M

Terminal Value after 2 years = 14* 47 = $ 658 M (since the exit multiple is 14)

Cost of capital = 11.2%

Enterprise value (EV) = 47/(1.12) + 47/(1.12)^2 + 658/(1.12)^2 = $ 603.98

MV of Equity = EV - Debt + Cash = 603.98 - 24 + 6 = $ 585.98 M

Stock price = (585.98/14) = 41.85 = $ 41.8 (rounded to one decimal)